the sum of three consecutive numbers, increased by 10, is equal to the sum of the first two consecutive number. find the numbers
Answer by reviewermath(1029)

x + (x+1) + (x+2) + 10 = x + (x+1)
3x + 13 = 2x + 1
x = -12
Answer: -12, -11, -10